White light reflected at a perpendicular incidence from a uniform soap film has an interference maximum at 666 nm, and a minimum at 555 nm, with no minima between 666 nm and 555 nm. If n=1.34 for the film, what is the film thickness?
1
Expert's answer
2020-05-18T10:06:20-0400

The condition for maximum in a thin film is:

"2dn+\dfrac{\lambda_1}{2}=m\lambda_1"

Where m is a whole number and do is the thickness of the film.

The condition for minimum will be:

"2dn+\dfrac{\lambda_2}{2}=(2m+1)\dfrac{\lambda_2} {2}"

Solving these two equations simultaneously, obtain:

"\dfrac{2dn} {\lambda_1} +0.5 =\dfrac{2dn} {\lambda_2}"

Thus:

"0.004d +0.5 =0.0048d\\
0.0008d=0.5\\
d=625 nm."


Answer. d=625 nm.